The capacitance of a parallel plate air capacitor is 350.0mF. When a sheet of a dielectric is inserted between the plates, the capacitance increases to 2100.0mF. What is the permittivity of the dielectric?

1
Expert's answer
2022-02-07T15:27:42-0500

For the empty and filled capacitor:


"C_e=\dfrac{\epsilon_0A}{d},\\\space\\
C_f=\dfrac{\epsilon\epsilon_0A}{d},\\\space\\
\dfrac{C_f}{C_e}=\epsilon,\\\space\\
\epsilon=2100/350=6."